OpenStudy (anonymous):

During a 4th of July weekend, 32 vehicles became trapped on the Sunshine Skyway Bridge while it was being repaved. A recent city ordinance decreed that only cars with 4 wheels and trucks with six wheels could be on the bridge at any given time. If there were 148 tires that needed to be replaced to due to damage, how many cars and trucks were involved in the incident?

OpenStudy (anonymous):

x=cars y= trucks

OpenStudy (anonymous):

32 cars so x=y+32

OpenStudy (anonymous):

car has 4 wheels so 4x and tucks have 6 wheels so 6x

OpenStudy (anonymous):

total number of wheels is 148 so 4x+6y=148

OpenStudy (anonymous):

x+y=32 4x+6y=148 so the answer is 148

OpenStudy (anonymous):

but how many cars?

OpenStudy (anonymous):

how many trucks?

OpenStudy (anonymous):

hold on let my think and type sorry

OpenStudy (anonymous):

mulptly the equation by-4 to get rid of x and then solve for y

OpenStudy (anonymous):

you said that 32 cars... its 32 vehicles... so would it be 32=x+y?

OpenStudy (anonymous):

-4x-4y=-128 4x+6y=148 2y=20 y=10 10 trucks and 22 cars where involved

OpenStudy (anonymous):

i dont understand ...

OpenStudy (anonymous):

4*{22}+10*6=88+60=148 thats to check it

OpenStudy (anonymous):

i dont know how to explin im sorry

OpenStudy (anonymous):

i just know how to do it... hold on...

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

does that help or no?

OpenStudy (anonymous):

Hey uselessmass ... want to help me with my question?? please!

OpenStudy (anonymous):

very much so. thanks

OpenStudy (anonymous):

np thats where i went when i had that problem.. ugh i had a whole work sheet of prob like ur and that website helped so i would say go there every time u dont get it that place

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np.... u do flvs ?

OpenStudy (anonymous):

yeah.. u?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what classes are you taking with flvs?

OpenStudy (anonymous):

what grade?

OpenStudy (anonymous):

i jus started and i only have two.. are u in flvs or pbv

OpenStudy (anonymous):

palm beach virtual = pbv?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

im in flvs... im taking 2 alg ii and bio i

OpenStudy (anonymous):

@Useless you can do it both ways right?

OpenStudy (anonymous):

im taking two too... and i have lang and alg honers in both it sucks

OpenStudy (anonymous):

alg what

OpenStudy (anonymous):

1 i think

OpenStudy (anonymous):

im in 9th

OpenStudy (anonymous):

u?

OpenStudy (anonymous):

im in 9th too!

OpenStudy (anonymous):

lol where u live in fl.. like ur city

OpenStudy (anonymous):

PBG u?

OpenStudy (anonymous):

lake worth city

OpenStudy (anonymous):

holly crap where did u go before taht?

OpenStudy (anonymous):

that*

OpenStudy (anonymous):

I was in the Jupiter schools...

OpenStudy (anonymous):

will do uselessmass

OpenStudy (anonymous):

i went to a charter and i got into some trouble ... kicked a back kids a ss cuz he touched my boob and decided i would rather be at home with my family and doing flvs on my own damn pace (fast)

OpenStudy (anonymous):

is that rly u in ur pic??

OpenStudy (anonymous):

i am hospital homebound... so it works both ways... one of my friends lives in lake worth... yeah that was taken two weeks ago

OpenStudy (anonymous):

ur a what>> wht is thtt sorry>??

OpenStudy (anonymous):

i was in and out of the hospital...i was missing so much school they put me on flvs... ive never returned to regular school

OpenStudy (anonymous):

oh pellet are you alright now?? also if u dont mind me asking why>?

OpenStudy (anonymous):

wait u r ok now? and why

OpenStudy (anonymous):

whos ur alg teac? mys is christy cole

OpenStudy (anonymous):

i have celiac disease

OpenStudy (anonymous):

mgtg... talk latter byee!

OpenStudy (anonymous):

mgtg?

OpenStudy (anonymous):

gtg*

OpenStudy (anonymous):

whats your email?

OpenStudy (anonymous):

it was nice talking to u .... ur cool iloveboys123@rocketmail.com email me anytime

OpenStudy (anonymous):

Thanks for everything

OpenStudy (anonymous):

anytime...whats ur name??

OpenStudy (anonymous):

ok i rly gtg or my papa is going to scream lol byee email! lol (:

OpenStudy (anonymous):

My name is Lars. I sent you an email. Reply :)